### v3

# CLUE: A Clinical Language Understanding Evaluation for LLMs

Abstract
Large Language Models (LLMs) are expected to significantly contribute to patient care, diagnostics, and administrative processes. Emerging biomedical LLMs aim to address healthcare-specific challenges, including privacy demands and computational constraints. Assessing the models’ suitability for this sensitive application area is of the utmost importance. However, evaluation has primarily been limited to non-clinical tasks, which do not reflect the complexity of practical clinical applications. To fill this gap, we present the Clinical Language Understanding Evaluation (CLUE), a benchmark tailored to evaluate LLMs on clinical tasks. CLUE includes six tasks to test the practical applicability of LLMs in complex healthcare settings. Our evaluation includes a total of $2525$ LLMs. In contrast to previous evaluations, CLUE shows a decrease in performance for nine out of twelve biomedical models. Our benchmark represents a step towards a standardized approach to evaluating and developing LLMs in healthcare to align future model development with the real-world needs of clinical application. We open-source all evaluation scripts and datasets for future research: https://github.com/TIO-IKIM/CLUE
CLUE: A Clinical Language Understanding Evaluation for LLMs
Amin Dada1   Osman Alperen Koraş 1   Marie Bauer1 Amanda Butler Contreras2 Kaleb E Smith2Constantin Marc Seibold1   Jens Kleesiek1,3,4,5 1Institute for AI in Medicine (IKIM), University Hospital Essen (AöR), Essen, Germany 2 NVIDIA, Santa Clara, CA, USA 3Cancer Research Center Cologne Essen (CCCE), West German Cancer Center Essen University Hospital Essen (AöR), Essen, Germany 4German Cancer Consortium (DKTK, Partner site Essen), Heidelberg, Germany 5Department of Physics, TU Dortmund, Dortmund, Germany
1 Introduction
Figure 1: The proposed CLUE Benchmark consists of six clinical tasks ranging from answering patient inquiries to multi-document knowledge extraction and simplification.
The introduction of Large Language Models (LLMs) into the healthcare sector marks the beginning of a transformative period for medical practitioners, promising significant advancements in the quality and efficiency of patient care Moor et al. (2023). In clinical research settings, these advances are driven by open-source LLMs, specifically the development of biomedical LLMs (Luo et al., 2023; Chen et al., 2023; Labrak et al., 2024). These models are tailored to the medical sector to deliver improved performance with fewer parameters than their general-domain counterparts.
Despite the promise these models hold, their evaluation has predominantly focused on constructed medical questions, such as exam quizzes (Han et al., 2023; Toma et al., 2023; Chen et al., 2023; Labrak et al., 2024; Gururajan et al., 2024; Griot et al., 2024). While indicative of the models’ medical knowledge, these assessments do not adequately reflect the complexity and diversity of clinical settings they are intended for.
Moreover, the number of released general and biomedical LLMs is immense and biomedical evaluations settings are often not overlapping or even contradict each other Touvron et al. (2023a); Han et al. (2023); Jiang et al. (2023); Touvron et al. (2023b); Toma et al. (2023); Chen et al. (2023); Li et al. (2023); Tunstall et al. (2023); Jiang et al. (2024); Labrak et al. (2024); Gururajan et al. (2024); Griot et al. (2024); Christophe et al. (2024); Abdin et al. (2024). This rapid evolution poses challenges in determining the most appropriate LLMs for clinical applications at any given moment, including assessing whether the latest general LLMs surpass the performance of previous domain-specific LLMs. To clarify this highly dynamic landscape, it is essential to implement ongoing evaluations using a standardized benchmark.
Catering to this demand, our study introduces the Clinical Language Understanding Evaluation (CLUE), a novel benchmark specifically designed to assess LLMs in the context of practical downstream tasks in healthcare in a comparable setting. CLUE is based on clinical documents and provides a comprehensive framework for evaluating the performance of LLMs. The benchmark consists of six clinical tasks that probe a diverse set of model abilities, including medical reasoning, information extraction, simplification, coding, and summarization, offering a robust challenge to test the essential capabilities of these models in a clinical setting. This allows us to analyze the strengths and shortcomings of $2525$ models. We include open-source models to address typical demands for data privacy regulations in clinical environments Liu et al. (2023); Wang et al. (2023).
In summary, our initiative to develop CLUE and conduct a thorough evaluation of both biomedical and general-domain LLMs aims to bridge the current knowledge gap in testing these technologies in healthcare. By challenging models with a wide set of clinical tasks, our benchmark seeks to pave the way for more informed decisions in the deployment of LLMs in medical settings, ultimately enhancing patient care and streamlining the workload of healthcare professionals.
2 Related Work
2.1 Large Language Models in the clinical domain
With the transformer architecture (Vaswani et al., 2017), automatic language processing gained noticeable traction (Hossain et al., 2019; Wolf et al., 2019), which has led to the development of initial foundation models for not only general language processing (Devlin et al., 2018; Conneau and Lample, 2019; Raffel et al., 2020) but also the specialized medical context (Liu et al., 2021; Yan et al., 2022; Rasmy et al., 2021). This progression has only been accelerated since the reveal of ChatGPT in late 2022 as the potential use for dynamically usable LLMs has been spread through the public. Commercial models such as Palm Chowdhery et al. (2022) and GPT4 (Achiam et al., 2023) seem to deliver great performances for various tasks that at first glance might rival specialists (Yunxiang et al., 2023; Nori et al., 2023; Lee et al., 2023; Thirunavukarasu et al., 2023), but on more thorough investigation show noticeable discrepancies (Lee et al., 2023). Shortly after the release of commercial models, a surge of open-source alternatives like Llama (Touvron et al., 2023a; Wu et al., 2023), Alpaca (Taori et al., 2023) and Mistral (Jiang et al., 2023) followed them. To deliver better performance on medical tasks, these models were pre-trained on medical data (Han et al., 2023; Peng et al., 2023a, b; Labrak et al., 2024). The specific evaluation settings, such as whether a model is fine-tuned on downstream tasks, how many examples are given, and techniques like chain-of-thought prompting, often vary, making the results impossible to compare. CLUE addresses this by evaluating models in a uniform setting on clinically relevant tasks.
2.2 Medical evaluation of LLMs
Figure 2: An overview of the MeDiSumQA generation process. In the first phase we use an LLM to split the discharge instructions to sentences. We then formulate question-answer pairs based on the sentence. These candidates are then manually check for our quality criteria. The hand-selected form the final MeDiSumQA dataset.
Evaluation of medical LLMs has predominantly focused on datasets comprising multiple choice exam questions, such as MMLU (Hendrycks et al., 2021), MEDQA (Jin et al., 2021), MedMCQA (Pal et al., 2022), and synthetic questions generated from PubMed articles (Jin et al., 2019). These tasks are a good way to test the representation of medical knowledge in models and, to some extent, their medical reasoning capabilities. However, they fail to address several critical aspects necessary for effective real-world medical applications, such as the comprehension of complex patient cases from semi-structured documents with medical jargon and abbreviations, the reformulation of information for tasks like summarization, the translation between lay and specialist language, and the answering of open-ended questions. Furthermore, recent work has reported severe issues with data contamination of LLM evaluation (Magar and Schwartz, 2022; Shi et al., 2024), specifically with regards to MMLU (Deng et al., 2023), and proposed solutions to mitigate these issues (Li et al., 2024; Deng et al., 2023).
This evaluation paradigm highlights a notable gap in that none of the models were assessed using clinical text datasets. Despite the requirement for extensive medical knowledge and reasoning in addressing these quizzes, the nature of the texts vastly differs from that of clinical documents. Clinical documents are characterized by their irregular structure, frequent jargon, typing errors, and abbreviations that vary across medical specialties and regions, and their typically greater length than quiz questions, leading to a less homogeneous domain of texts.
DR.BENCH (Gao et al., 2023) is a benchmark focusing on clinical documents but predominantly consists of traditional encoder model tasks, such as extractive question-answering, sequence tagging, and text classification, rendering it less suitable for a comprehensive evaluation of LLMs. Kweon et al. (2024) introduced a multiple-choice question-answering dataset derived from MIMIC-IV Electronic Health Records (EHR), and Adams et al. (2024) released another MCQA dataset based on fictive reports concerning constructed patient cases. However, both studies refrained from evaluating biomedical LLMs on their datasets. CLUE combines appropriate existing tasks with novel tasks for a standardized clinical evaluation of general-domain and biomedical models.
3 CLUE Benchmark
CLUE includes six clinical document tasks, which encompass information extraction, summarization, clinical reasoning, simplification, and coding. These tasks are designed to better assess the practical applications of LLMs in clinical settings. Table 1 in Appendix summarizes the characteristics of these tasks. We provide prompt examples for each task in Figures 9, 10, 11, 12, 13, and 14 in Appendix B.3.
3.1 MeDiSumQA
MeDiSumQA is a novel medical open question-answering dataset we derived from MIMIC-IV (Johnson et al., 2021) discharge summaries.
3.1.1 Dataset Generation
Some of the discharge summaries contain a discharge letter to the patient. This letter summarizes the central information of the report in condensed and simple language and contains follow-up instructions. Based on these letters, we generate question-answer pairs using Meta’s Llama-3-70B-Instruct AI@Meta (2024). The questions are formulated from the patient’s view, and the answers are based on information that was given to the patient. For construction, we filter for documents containing a specific string111“You were admitted to the hospital” that marks the beginning of the discharge letter. The creation process is divided into three phases that are visualized in Figure 2.
In the first phase, we split the text into sentences (for details, see A.1). In the second phase, we prompted the LLM to generate question-answer pairs from the list of sentences for each document. We instruct the LLM to ground each answer in one or more sentences so we can check the quality manually. After completing the first two phases, we have a list of question-answer pairs. In the third phase, these candidates are manually reviewed, and high-quality examples are selected based on the following criteria:
  1. Detail: We observed that some information in discharge letters did not match the rest of the discharge summary due to medical details that were not communicated in the letter.
  2. Difficulty: We found that some generated answers were too obvious, or the questions were so specific that they gave away the answer.
  3. Ambiguity: Many questions lacked clear answers. For instance, a question about a procedure performed for diagnosis is answered by mentioning a CT scan. However, the full report also listed laboratory tests as diagnostically relevant.
Figure 8 in Appendix A.1 lists some examples of problematic pairs and selected pairs.
As a last step, we remove the discharge letters from the original documents to pair the questions and answers with the resulting documents.
3.1.2 Task Qualities
This task requires models to perform multiple skills simultaneously. They must read and comprehend discharge summaries to fully understand the patient’s case, extract relevant details to answer specific questions about the patient’s condition and treatment and rephrase and simplify the information to ensure it is easily understandable by the patient. Additionally, the models need a robust understanding of medical knowledge and clinical guidelines to infer and provide appropriate follow-up advice accurately.
This multifaceted challenge assesses the LLM’s ability to integrate comprehension, extraction, communication, and domain-specific expertise in a clinical context.
3.2 MeDiSumCode
Figure 3: An overview of the MeDiSumCode dataset.
Coding of discharge summaries involves assigning International Classification of Diseases (ICD) (Organization, 2004) codes to diagnoses and procedures. This task is necessary for patient records, billing, and statistical analysis in healthcare.
3.2.1 Dataset Generation
Using MIMIC-IV, we generate an ICD-10 prediction dataset by leveraging the hosp module, which contains manually annotated ICD-10 codes for primary and secondary diagnoses and corresponding discharge summaries. Figure 3 shows that our process links discharge summaries to their respective ICD-10 codes using the unique hospital admission ID. This linkage allows us to create a cohesive dataset where each discharge summary is accurately paired with its diagnostic codes, providing both the inputs (discharge summaries) and the labels (ICD-10 codes) for model evaluation.
3.2.2 Task Qualities
This task is challenging for several reasons. First, it requires accurate identification of relevant diagnoses from complex and long clinical text. Second, the system must have detailed knowledge of the vast ICD-10 coding system, which contains over 70,000 codes, in order to accurately match these diagnoses. This requires a large database of codes and an understanding of the medical conditions they represent. Finally, matching diagnoses to the correct ICD-10 codes involves reasoning to combine both prior steps for a prediction.
MeDiSumCode assesses an LLM’s ability to extract relevant information from complex text accurately, possess detailed knowledge of a specialized coding system, and employ reasoning to integrate these elements for precise predictions.
3.3 MedNLI
The basis for MedNLI (Romanov and Shivade, 2018) are clinical notes from the MIMIC III (Johnson et al., 2016) dataset. Sentences are sampled from medical history sections and given to clinicians. Based on each sentence, generate three hypotheses—contradictory, neutral, and entailed—. Models are tested by predicting the logical relationship between the premise and these hypotheses.
The input length of this dataset is extremely small, making it well suited to measuring the ability for clinical reasoning regardless of the difficulty of LLMs on long inputs.
3.4 MeQSum
MeQSum (Ben Abacha and Demner-Fushman, 2019) consists of 1,00010001,0001 , 000 consumer health inquiries from the U.S. National Library of Medicine that medical experts manually summarized. This task provides a good complement to clinical documents written by medical staff, as it checks whether models are able to understand lay language.
MeQSum probes if LLMs are able to identify key information patient queries and reformulate them to brief, directed medically sound queries.
3.5 ProblemSummary
This task was first described by Gao et al. (2022) and utilizes clinical notes organized according to the widely recognized SOAP principle, which divides notes into Subjective, Objective, Assessment, and Plan sections (Weed, 1964). For this task, only the Subjective and Assessment sections were employed as inputs for the models, with the Objective portion being excluded. Based on this information, models are assigned to predict a patient’s current health problems.
Similar to MedNLI, the input length of this dataset is quite short, allowing us to test basic information extraction abilities without requiring a long context window.
3.6 LongHealth
The LongHealth (Adams et al., 2024) dataset contains 20 fictional patient records with various diseases. The reports are designed to be long and thus present an additional challenge to LLMs, as their performance has previously been shown to worsen with increasing input length (Levy et al., 2024). Evaluation is split into three sub-tasks: The first sub-task measures how well models can answer questions about long documents. Sub-task two increases the input length by including unrelated documents. The third sub-task removes the relevant document, and a new answer option is added to state that the information is not available.
This task measures a model’s comprehension of complex clinical documents, its ability to retain this performance on long inputs, and its tendency to hallucinate an answer when information is not available.
3.7 Metrics
For open-ended tasks, we report the F1-score between the model predictions and ground truth unigrams (ROUGE-1), bigram (ROUGE-2), and the longest common subsequence (ROUGE-L) [^2] https://huggingface.co/spaces/evaluate-metric/rouge (Lin, 2004). We compute the BERTScore (Zhang et al., 2019) on clinical documents to measure semantic similarity using an encoder trained on MIMIC III333emilyalsentzer/Bio_ClinicalBERT (Alsentzer et al., 2019). We first tuned the score rescaling baselines for MIMIC IV discharge summaries. For Problem Summaries and MeDiSumQA, we also extract the Unified Medical Language System (UMLS) (Bodenreider, 2004) entities with scispacy (Neumann et al., 2019) and compute their F1-score to consider medical abbreviations and synonyms. When evaluating MedDiSumCode, we calculate the ratio of valid ICD-10 codes. We use the python package icd10-cm [^4] https://pypi.org/project/icd10-cm/ to probe the validity of ICD-10 codes. We distinguish between exact match (EM) and the match of the first three characters of the codes, which is an approximate match (AP) based on the hierarchical structure of ICD-10 codes.
Following the SuperGLUE benchmark (Wang et al., 2019), we compute an aggregate score for both levels by averaging all over all datasets. For datasets with more than one metric, we first compute the average score of the dataset.
4 Experimental setup
Figure 4: The average level 1 and level 2 scores. Models that share the same base are displayed in the same color.
We evaluated $2525$ language models, including general domain and specialized biomedical models. Since some biomedical models were not instruction-tuned, we also included general domain base models for a fair comparison. Our evaluation aims to (1) rank current openly available models, (2) assess whether biomedical models or general domain instruction-tuned models are more suitable for specific clinical scenarios, and (3) measure the effects of continuous biomedical training.
4.1 General-domain models
We include both the instruction-tuned and base checkpoints of Llama2-7B, Llama2-70B Touvron et al. (2023b), Mistral-7B-v0.1 Jiang et al. (2023), Meta-Llama-3-8B-Instruct and Meta-Llama-3-70B-Instruct AI@Meta (2024) since they serve as a basis for several biomedically fine-tuned models. Additionally, we include zephyr-7b-beta Tunstall et al. (2023), Mistral-7B-Instruct-v0.2 Jiang et al. (2023), Phi-3-mini-128k-instruct Abdin et al. (2024), Mixtral-8x7B and Mixtral-8x22B Jiang et al. (2024) due to their general popularity.
Since the context length of Llama2-based models is insufficient to read all MIMIC discharge summaries and LongHealth documents, we split the evaluation into two levels: The first one with input sizes of less than 1.000 words on average and level two with more than 1.000 on average with long documents. Table 1 in Appendix A shows the average number of words per input for all tasks.
4.2 Biomedical Models
Table 2 in B.2 shows the biomedical LLMs we evaluate in our experiments. Meditron-7B and 70B (Chen et al., 2023) are continuously pretrained from Llama-2-7B Touvron et al. (2023a). The pretraining dataset consists of 48.1B tokens (Chen et al., 2023). In contrast to Meditron, internistai/base-7b-v0.2 (Griot et al., 2024) was only trained on 2.3B tokens with a focus on data quality based on Mistral-7B-v0.1 (Jiang et al., 2023). BioMistral (Labrak et al., 2024) is another continuously pre-trained model that was based on Mistral-7B-Instruct-v0.1. They use weight-merging algorithms such as DARE (Yu et al., 2023) to combine the weights of their biomedical model and the original instruction-tuned model. Toma et al. (2023) released ClinicalCamel, a language model based on Llama-2-70B. A more recent model was published by Gururajan et al. (2024) and is based on Meta’s Llama3 (AI@Meta, 2024). They employed Direct Preference Optimization (DPO) and supervised fine-tuning (SFT) to train instruction-tuned models. A similar fine-tuning procedure was done for Llama3-OpenBioLLM-8B and Llama3-OpenBioLLM-70B (Ankit Pal, 2024). However, no information on the training process has been provided yet. Finally, the models JSL-Med-Sft-Llama-3-8B John Snow Labs (2024a), JSL-MedLlama-3-8B-v1.0 John Snow Labs (2024b) and JSL-MedLlama-3-8B-v2.0 John Snow Labs (2024c) also fine-tuned Llama3 variants with very limited available information.
Figure 5: The relative difference between biomedical models and their base model.
4.3 Baselines
For each score, we provide a lower bound for meaningful predictions as a baseline to facilitate the interpretability of the scores. We use the expected value for random answer selection for the two data sets with finitely many answers, MedNLI and LongHealth. For MeQSum and MeDiSumQA, we use the unchanged input as a prediction. In the case of Problem Summary, we extract all UMLS entities from the input and use them as predictions. Finally, for MeDiSumCode, we collect all ICD-10 codes occurring in the dataset and draw random codes from this set.
5 Results
We present the benchmark results in Figures 4 and 5. Detailed scores for all models are in Table 3 of Appendix C.
Figure 4 shows that overall the general-domain models Meta-Llama-3-70B-Instruct, Mixtral-8x22B-Instruct-0.1 and Meta-Llama-3-8B-Instruct score the highest. The biomedical model Llama3-OpenBioLLM-70B is also among these models, however clearly behind its base model Meta-Llama-3-70B-Instruct. The lower left quadrant is mainly filled with biomedical models, with some scoring below the baseline.
For well-performing models, the difference between level 1 scores is generally smaller than the difference between level 2 scores. This indicates the expressiveness of LongHealth and the newly introduced tasks MeDiSumQA and MeDiSumCode.
5.1 Effect of biomedical fine-tuning
For many biomedical models, there is a clear decline in performance, which is visualized in Figure 5. While models based on Mistral-7B-v0.1 improve up to $5.525.52$ points, models based on Meta-Llama-3 show a very clear decrease with drops of up to $31.3731.37$ points. The remaining models, Meditron-7B, Meditron-70B, and ClinicalCamel-70B, are also not able to improve the original Llama2-70B performance, while the instruction-tuned general domain Llama2-70B very clearly outperforms the biomedical models. (see Appendix C).
5.2 Error Analysis
Major contributions to the performance drop of biomedical models are task 3 of LongHealth and the MeDiSumCode valid code scores. Since LongHealth Task 3 measures how often a model returns no answer if the requested information is not in the input document, it indicates the frequency of hallucinations. Similarly, the valid code scores in MeDiSumCode show models often invent ICD-10 codes. Models that have a low valid code score often start to count up the numbers in ICD codes instead of actually predicting valid codes. Examples of this behavior are listed in Appendix C.1. LongHealth task 3 reveals a drastic change from Meta-Llama-3-8B-Instruct, which scored $56.2556.25$, to Llama3-OpenBioLLM-8B, which only scored $1.551.55$. Similarly, Llama3-OpenBioLLM-70B also performs poorly compared to Meta-Llama-3-70B-Instruct on this task.
In addition to increased hallucinations, we observed that biomedically fine-tuned models often exhibit generation loops, wherein they repetitively produce the same sequences without reaching a conclusion. This looping behavior significantly hampers the models’ ability to generate coherent and relevant responses. Examples of this phenomenon have been observed in various tasks, highlighting the consistent nature of these generation loops.
Another common issue identified in poorly performing models is their inability to adhere to the prescribed instruction format, particularly in tasks involving long inputs at level 2. This problem is notably pronounced in tasks such as LongHealth, where the expectation is for models to produce answers in a strict, parseable format.
6 Discussion
6.1 Effect of biomedical training
Our analysis reveals a notable decline in performance for several models following biomedical training, observable after both continuous pre-training and methods like SFT and DPO. An outlier in this trend is BioMistral-DARE, which employed weight merging with the original instruction-tuned checkpoint. This suggests that such a training approach may mitigate the adverse effects of fine-tuning.
However, the superior performance of Mistral-7B-Instruct-v0.2, which shares the same architecture as BioMistral-7B-DARE, implies that an enhanced general domain pre-training dataset might have a more profound impact than fine-tuning alone. Assessing this conclusively is challenging, given the lack of details regarding the pre-training data for Mistral-7B-Instruct-v0.2.
Additionally, many models trained using SFT relied on generated data, hinting that performance issues may stem from data quality. internistai/base-7b-v0.2, which was trained on high-quality data, showed the highest improvement in performance on level 2 tasks, supporting the hypothesis that data quality is crucial.
All biomedical models that showed improvement are based on the relatively low-scoring Mistral-7B-Instruct-v0.1. This leads us to our final hypothesis: these improvements might address gaps that more recent general-domain models, such as Meta-Llama-3 and Mistral-7B-Instruct-v0.2, have already overcome.
Therefore, we conclude that biomedical training with currently existing methods and datasets is not beneficial for models in clinical settings. Future research could be focused on novel training methods that counteract model performance deterioration or focus on datasets that address this issue.
6.2 Concerns About Data Contamination
Recently, data contamination has emerged as a significant concern in LLMs, particularly for open benchmarks like Huggingface’s Open LLM Leaderboard. This problem arises when the datasets used for evaluation have overlaps with training datasets, or when LLMs have been pre-trained or fine-tuned on the evaluation datasets themselves. Such contamination compromises the integrity of the evaluation process, skewing test results by favoring models for their memorization capabilities rather than their ability to generalize. Consequently, this has spurred a surge in research to address this issue (Li et al., 2024; Deng et al., 2023).
Accidental data contamination is less likely in datasets derived from sources such as MIMIC, where, for example, ICD-10 codes are not directly available but must be extracted separately. However, the risk of intentional benchmark manipulation persists. One approach to mitigate this is through the involvement of trusted entities, such as hospitals that have access to private data, to create their own confidential datasets. To foster the development of new datasets based on report summaries and allow users to assess LLMs with their private data, we will make our source code for the dataset generation process publicly available.
7 Conclusion
In conclusion, our study reveals that current biomedical LLMs are not competing effectively with general-domain models in clinical settings. While some biomedical models improved marginally, more recent models based on Meta-Llama-3 are underperforming. Fine-tuning these models with domain-specific data often results in decreased performance, introducing issues such as hallucinations and reduced model stability. This starkly contrasts traditional MCQA evaluations, such as USMLE, where biomedical models have previously shown superior performance. The introduction of the CLUE benchmark highlights these discrepancies, providing a more accurate and practical assessment of LLM capabilities in real-world healthcare applications. These findings underscore the need for ongoing evaluation and refinement to ensure the effective deployment of LLMs in medical practice.
Limitations
Our study has several limitations that should be considered. Due to the significant computational resources required to run LLMs with up to 141 billion parameters, we did not explore the impact of various model configurations, such as temperature settings, or advanced techniques like chain-of-thought prompting on model performance. Future research should investigate these aspects to gain a more comprehensive understanding of their effects. Additionally, our new datasets are based on publicly available data from MIMIC-IV. As such, we cannot completely prevent data contamination. This limitation underscores the need for future research into robust methods for mitigating data contamination, which is crucial for ensuring the validity of any public LLM benchmark. While we presented novel insights in this paper, their application to clinical data requires further investigation. Future work should refine these methods to enhance their applicability and reliability in clinical settings. Furthermore, our evaluation primarily focused on tasks involving clinical documents and their relevance, but it was not conducted in a realistic clinical setting. Therefore, extensive evaluation through prospective clinical trials is necessary to meet the required safety levels before applying these models to clinical environments.
Appendix A Task Details
The characteristics of all included tasks are summarized in Table 1.
Dataset Samples Mean Words Metrics Focus
MedNLI 1425 20.81 Accuracy Clinical reasoning
MeQSum 1000 60.77 R-L, R-1, R-2, BERT F1 Summarization
LongHealth 400 5536.82 Accuracy Information extraction
Prob.Sum. 237 123.5 R-L, R-1, R-2, BERT F1, UMLS F1 Information extraction
MeDiSumQA 453 1451.79 R-L, R-1, R-2, BERT F1, UMLS F1 Simplification/Clinical reasoning
MeDiSumCode 500 1515.32 EM F1, AP F1, Valid Code Information extraction / Coding
Table 1: An overview of the characteristics of the tasks CLUE consists of.
A.1 MeDiSumQA Generation
Figure 6: The prompt we use to split the discharge instructions into sentences.
Figure 7: The prompts we are using to generate MeDiSumQA.
We do not use a classic sentence splitter (e.g., NLTK), as these do not work reliably due to the irregular formatting and placeholders introduced by anonymization. Instead, we prompt an LLM to split the text into sentences without changing anything else. We make sure that the LLM has not changed the sentences by checking whether the sentences can be found in the original document via exact string matching. Figure 6 shows our prompt for this.
Figure 8: This figure shows representative examples of issues we detected with some of the generated questions-answer pairs.
We manually select the final examples for the dataset based on correctness, difficulty, and ambiguity. Figure 8 shows representative examples of pairs we filtered out.
Model Name Base Model Type of Training
Mditron-7b llama2-7b Continued pretraining
Internist.ai 7b Mistral-7B-v0.1 Continued pretraining + SFT
BioMistral-7B Mistral-7B-Instruct-v0.1 Continued pretraining
BioMistral-7B-DARE Mistral-7B-Instruct-v0.1 Continued pretraining +DARE
Llama3-OpenBioLLM-8B Meta-Llama-3-8B-Instruct SFT + DPO
JSL-Med-Sft-Llama-3-8B Meta-Llama-3-8B-Instruct Unknown
JSL-MedLlama-3-8B-v1.0 Meta-Llama-3-8B-Instruct Unknown
JSL-MedLlama-3-8B-v2.0 Meta-Llama-3-8B-Instruct Unknown
Llama3-Aloe-8B-Alpha Meta-Llama-3-8B-Instruct SFT + DPO
Meditron-70B Llama-2-70b Continued pretraining
ClinicalCamel-70B llama2-70b SFT
Llama3-OpenBioLLM-70B Meta-Llama-3-70B-Instruct SFT + DPO
Table 2: Evaluated Biomedical Models
Appendix B Experimental setup
B.1 Computational Resources
All experiments were conducted on an NVIDIA DGX A100 640GB node with 8x NVIDIA A100 80GB Tensor Core GPUs within three days, resulting in approximately 1536 GPU hours.
B.2 Models
Table 2 lists all biomedical models we evaluated.
B.3 Prompting
We apply few-shot prompting and use the instruction template on Hugging Face for the instruction-tuned models. For the other models, we concatenate the system prompt, few-shot examples, and user prompt into one string separated by double newlines. For the level one evaluation, we performed 3-shot prompting. For level two, we provide one shot with the exception of LongHealth, where we provide no examples due to the content length.
Figures 9, 10, 11, 12, 13, and 14 are showing the prompt formats we are using for the different benchmark tasks. If the input length allowed this, we also included sample texts from the datasets.
Figure 9: MeQSum prompt format with example.
Figure 10: Problem Summary prompt format with example.
Figure 11: MedNLI prompt format with example.
Figure 12: LongHealth prompt format.
Figure 13: MeDiSumQA prompt format.
Figure 14: MeDiSumCode prompt format.
Appendix C Results
Table 3 shows the detailed benchmark results for all models.
Figure 15: Biomedical models that show the described counting behavior compared to their base model.
C.1 Error Analysis
Figure 15 shows some examples of the described type of error with regard to counting.
MedNLI Problem Summary MeQSum
Level 1 Score Acc R-L R-1 R-2 BERT F1 UMLS F1 R-L R-1 R-2 BERT F1
baseline 25.13 33.33 6.22 8.53 1.58 61.07 8.24 14.96 18.99 7.21 58.62
llama2-7b Touvron et al. (2023a) 20.11 29.51 5.97 7.35 2.11 59.45 9.06 7.16 8.58 2.89 37.5
meditron-7b Chen et al. (2023) 11.46 2.39 6.49 7.87 2.59 59.57 12.52 6.17 7.47 2.53 40.53
llama2-7b-hf Touvron et al. (2023a) 36.94 41.27 17.43 22.25 6.93 66.33 21.62 36.45 39.99 18.13 75.98
Mistral-7B-v0.1 Jiang et al. (2023) 33.99 67.54 7.07 8.99 3.17 60.74 15.67 6.28 7.86 2.49 44.6
internistai/base-7b-v0.2 Griot et al. (2024) 25.98 50.00 4.67 6.08 1.59 57.87 6.8 5 5.85 0.74 38.54
BioMistral-7B Labrak et al. (2024) 40.79 62.32 18 21.85 8.63 66.96 20.92 24.58 27.38 12.1 67.05
BioMistral-7B-DARE Labrak et al. (2024) 43.24 66.76 18.81 22.92 8.93 68.93 22.65 26.16 29.69 13.49 68.68
Mistral-7B-Instruct-v0.1 Jiang et al. (2023) 40.31 64.79 14.46 18.6 6.26 65.79 20.08 21.85 25.26 11.17 66.15
zephyr-7b-beta Tunstall et al. (2023) 42.69 68.45 14.81 19.91 5.98 67.39 19.2 25.66 29.81 12.33 68.85
Mistral-7B-Instruct-v0.2 46.45 69.93 19.57 25.59 8.92 69.64 22.07 33.54 37.47 16.61 73.47
Meta-Llama-3-8B-Instruct 48.37 74.08 22.7 28.52 9.87 71.45 25.32 32.2 36.49 16.37 72.74
Llama3-OpenBioLLM-8B 33.20 44.93 10.82 13.62 4.03 64.14 15.67 26.21 29.41 14.03 62.39
Llama3-Aloe-8B-Alpha Gururajan et al. (2024) 42.50 73.94 9.47 12.3 4.06 65.56 15 23.78 27.02 12.31 66.01
JSL-Med-Sft-Llama-3-8B 19.00 29.08 4.59 5.75 1.61 57.69 8.75 6.57 7.73 2.99 31.70
JSL-MedLlama-3-8B-v1.0 17.00 13.31 12.87 15.40 4.73 65.18 18.42 6.02 7.19 2.43 41.78
JSL-MedLlama-3-8B-v2.0 31.26 35.70 8.16 10.61 2.94 64.48 13.17 30.93 34.54 16.31 71.03
Phi-3-mini-128k-instruct 41.70 57.25 19.8 23.72 8.47 70.27 21.06 32.05 35.85 15.81 73.01
Mixtral-8x7B-Instruct-v0.1 Jiang et al. (2024) 47.81 76.48 17.44 23.39 7.7 68.56 19.51 32.47 36.38 16.86 72.8
llama2-70b Touvron et al. (2023b) 35.17 76.27 7.13 8.77 3.15 59.61 14.34 3.75 4.15 1.05 33.59
meditron-70b Chen et al. (2023) 30.58 63.52 7.22 8.91 3.21 60.08 13.91 2.95 3.24 0.76 31.27
ClinicalCamel-70B Toma et al. (2023) 35.71 64.65 8.62 10.83 3.71 59.94 12.34 16.96 18.8 8.49 49.3
llama2-70b-hf Touvron et al. (2023b) 42.90 61.69 14.43 19.81 6.14 65.07 21.37 34.91 38.81 18.48 74.37
Meta-Llama-3-70B-Instruct 52.36 79.37 25.43 33.16 13.01 73 29.12 36.57 40.2 19.3 75.74
Llama3-OpenBioLLM-70B 47.57 80.85 12.1 16.67 5.58 66.51 17.74 30.72 34.31 15.55 71.99
Mixtral-8x22B-Instruct-v0.1 51.88 84.93 17.97 22.6 8.17 69.29 22.31 36.15 39.94 19.45 75.02
LongHealth MeDiSumQA MeDiSumCode
Level 2 Score Task 1 Task 2 Task 3 R-L R-1 R-2 BERT F1 UMLS F1 EM F1 AP F1 Valid Code
baseline 24.86 20.00 20.00 16.66 13.11 15.76 2.82 60.22 12.74 0.88 3.44 100.00
Mistral-7B-v0.1 9.55 0.60 0.55 0.15 5.93 7.16 1.59 53.47 7.47 0.77 5.32 33.21
internistai/base-7b-v0.2 22.01 46.10 25.40 24.55 9.1 11.77 3.23 61.64 12.54 1.18 6.06 35.84
BioMistral-7B 12.23 5.20 2.55 0.90 14.65 17.81 5.46 59.01 16.88 0.65 4.57 27.88
BioMistral-7B-DARE 25.82 46.00 40.85 4.6 17.01 20.87 6.92 65.17 18.45 1.20 6.66 56.04
Mistral-7B-Instruct-v0.1 23.12 45.75 40.65 3.65 16.62 21.34 6.95 65.68 16.77 0.57 3.78 37.25
zephyr-7b-beta 28.17 42.9 30.5 26.35 13.02 17.68 4.98 64.08 13.92 2.31 12.00 71.27
Mistral-7B-Instruct-v0.2 38.94 67.2 62.4 42.45 21.8 27.47 9.19 68.43 20.3 3.08 18.23 68.76
Meta-Llama-3-8B-Instruct 40.61 68.3 66.55 56.25 22.44 28.15 9.63 68.62 22.74 3.95 17.55 61.93
Llama3-OpenBioLLM-8B 25.44 37.55 41.75 1.55 22.89 27.95 10.45 68.7 22.15 0.84 4.84 51.16
Llama3-Aloe-8B-Alpha 30.31 66.75 63.3 19.05 11.03 14.97 4.57 63.87 12.68 1.77 12.78 44.84
JSL-Med-Sft-Llama-3-8B 15.68 0.00 0.00 0.00 8.85 11.74 3.21 58.38 10.54 3.09 14.03 68.37
JSL-MedLlama-3-8B-v1.0 29.94 52.85 52.05 0.05 19.06 22.7 7.63 67.29 20.83 1.87 12.80 67.30
JSL-MedLlama-3-8B-v2.0 18.02 1.10 0.30 0.05 19.13 24.15 8.56 67.41 20.49 1.41 11.13 64.31
Phi-3-mini-128k-instruct 24.39 27.25 23.9 0.00 18.47 22.45 6.94 65.72 16.52 0.43 3.63 86.24
Mixtral-8x7B-Instruct-v0.1 42.57 76.5 73.65 24.2 20.72 26.4 8.96 67.74 20.25 10.49 28.99 82.87
Meta-Llama-3-70B-Instruct 56.00 81.65 77.90 91.70 26.2 32.5 11.93 70.24 25.78 19.65 39.2 93.94
Llama3-OpenBioLLM-70B 45.56 80.2 75.60 62.90 21.85 27.8 9.54 68.43 22.47 7.37 20.24 73.65
Mixtral-8x22B-Instruct-v0.1 51.23 79.00 73.90 86.30 21.81 27.75 9.42 68.51 22.70 15.95 35.96 79.84
Table 3: The benchmark results for level 1 and level 2. Models with the same base model are grouped. †Indicates biomedical training. The relative level of the scores for each column is colored from red for low scores to green for high scores.