### v4

# Does Biomedical Training Lead to Better Medical Performance?

Abstract
Large Language Models (LLMs) are expected to significantly contribute to patient care, diagnostics, and administrative processes. Emerging biomedical LLMs aim to address healthcare-specific challenges, including privacy demands and computational constraints. Assessing the models’ suitability for this sensitive application area is of the utmost importance. However, biomedical training has not been systematically evaluated on medical tasks. This study investigates the effect of biomedical training in the context of six practical medical tasks evaluating $2525$ models. In contrast to previous evaluations, our results reveal a performance decline in nine out of twelve biomedical models after fine-tuning, particularly on tasks involving hallucinations, ICD10 coding, and instruction adherence. General-domain models like Meta-Llama-3.1-70B-Instruct outperformed their biomedical counterparts, indicating a trade-off between domain-specific fine-tuning and general medical task performance. We open-source all evaluation scripts and datasets to support further research in this critical area.
Does Biomedical Training Lead to Better Medical Performance?
Amin Dada1   Osman Alperen Koraş 1   Marie Bauer1 Amanda Butler Contreras2 Kaleb E Smith2Constantin Marc Seibold1   Jens Kleesiek1,3,4,5 1Institute for AI in Medicine (IKIM), University Hospital Essen (AöR), Essen, Germany 2 NVIDIA, Santa Clara, CA, USA 3Cancer Research Center Cologne Essen (CCCE), West German Cancer Center Essen University Hospital Essen (AöR), Essen, Germany 4German Cancer Consortium (DKTK, Partner site Essen), Heidelberg, Germany 5Department of Physics, TU Dortmund, Dortmund, Germany
1 Introduction
Figure 1: Our evaluation consists of six medical tasks with increasing complexity ranging from answering patient inquiries to multi-document knowledge extraction and simplification.
The introduction of Large Language Models (LLMs) into the healthcare sector marks the beginning of a transformative period for medical practitioners, promising significant advancements in the quality and efficiency of patient care Moor et al. (2023). In medicine, a major factor for these advances are open-source LLMs, specifically biomedical LLMs (Luo et al., 2023; Chen et al., 2023; Labrak et al., 2024). These models are tailored to the medical sector to deliver improved performance with fewer parameters than their general-domain counterparts.
Despite the promise these models hold, their evaluation has predominantly focused on constructed medical questions, such as exam quizzes (Han et al., 2023; Toma et al., 2023; Chen et al., 2023; Labrak et al., 2024; Gururajan et al., 2024; Griot et al., 2024). While multiple-choice quizzes offer a controlled environment for assessment, real-world medical tasks are often open-ended, involve incomplete information that the model must recognize to avoid generating inaccurate responses, and require processing lengthy, unstructured inputs that challenge the model’s ability to organize and extract relevant data efficiently. This raises the question of whether biomedical training methods are also improving LLM performance on more practical medical tasks.
Dataset Samples Mean Words Metrics Focus
MedNLI 1425 20.81 Accuracy Clinical reasoning
MeQSum 1000 60.77 R-L, R-1, R-2, BERT F1 Summarization
LongHealth 400 5536.82 Accuracy Information extraction
Problem Summary 237 123.5 R-L, R-1, R-2, BERT F1, UMLS F1 Information extraction
MeDiSumQA 453 1451.79 R-L, R-1, R-2, BERT F1, UMLS F1 Simplification/Clinical reasoning
MeDiSumCode 500 1515.32 EM F1, AP F1, Valid Code Information extraction / Coding
Table 1: An overview of the characteristics of the tasks.
Moreover, the number of released general and biomedical LLMs is immense and biomedical evaluations settings are often not overlapping or even contradict each other Touvron et al. (2023a); Han et al. (2023); Jiang et al. (2023); Touvron et al. (2023b); Toma et al. (2023); Chen et al. (2023); Li et al. (2023); Tunstall et al. (2023); Jiang et al. (2024); Labrak et al. (2024); Gururajan et al. (2024); Griot et al. (2024); Christophe et al. (2024b); Abdin et al. (2024). Different settings are often used for evaluation, which makes it even more difficult to assess the general effects of biomedical training of models.
Our study addresses these challenges by evaluating LLMs before and after biomedical training on a set of practical downstream tasks in healthcare in a comparable setting. We evaluate on six tasks that probe a diverse set of model abilities, including medical reasoning, information extraction, simplification, ICD coding, and summarization, offering a robust challenge to test a range of essential capabilities of these models in a medical setting. Figure 1 summarizes the key features of these tasks. This allows us to analyze the strengths and shortcomings of $1212$ models before and after biomedical training.
Despite biomedically trained models often outperforming their general domain counterparts on multiple-choice quizzes, our study concludes that these results do not always translate to practical tasks. In fact, multiple biomedical LLMs perform worse on practical tasks than general models. This can be attributed to increased hallucinations and reduced accuracy and stability, suggesting that while specialized training enhances domain-specific knowledge, it may also introduce overfitting or bias, leading to unreliable outputs in complex, real-world medical scenarios. By challenging models with a wide set of medical tasks, our evaluation seeks to pave the way for more informed decisions in the deployment of LLMs in medical settings, ultimately enhancing patient care and streamlining the workload of healthcare professionals.
2 Related Work
2.1 Large Language Models in the clinical domain
The development of biomedical Large Language Models (LLMs) has gained considerable momentum in recent years, driven by the need for more specialized tools in the healthcare sector. This rapid growth is evident in the emergence of both commercial models, such as Med-PaLM (Singhal et al., 2023) and MedGemini (Saab et al., 2024), and open-source alternatives like Meditron (Chen et al., 2023), Biomistral (Labrak et al., 2024), Internist.ai (Griot et al., 2024), and Med42 (Christophe et al., 2024b).
Biomedical LLMs have shown a marked improvement over their general-domain counterparts, particularly in multiple-choice question-answering (MCQA) tasks, where they benefit from their domain-specific training. In some cases, these models have even challenged the performance of leading commercial tools like GPT-4 (Christophe et al., 2024a), highlighting their potential in handling specialized medical knowledge.
However, while biomedical fine-tuning has led to improved performance on MCQA tasks, it remains unclear whether these gains translate to practical medical tasks.
2.2 Medical evaluation of LLMs
Figure 2: An overview of the MeDiSumQA generation process. In the first phase we use an LLM to split the discharge instructions to sentences. We then formulate question-answer pairs based on the sentence. These candidates are then manually check for our quality criteria. The hand-selected form the final MeDiSumQA dataset.
Evaluation of medical LLMs has primarily centered on datasets composed of multiple-choice questions, such as MMLU (Hendrycks et al., 2021), MEDQA (Jin et al., 2021), MedMCQA (Pal et al., 2022), and synthetic questions from PubMed articles (Jin et al., 2019). While these tasks assess medical knowledge representation and some reasoning abilities, they overlook crucial aspects needed for real-world applications, such as understanding complex patient cases, translating between lay and specialist language, and addressing open-ended questions. This evaluation approach reveals a gap—models aren’t tested on clinical text datasets. Unlike exam questions, clinical documents are irregularly structured, rife with jargon, errors, abbreviations, and generally longer, making them a more heterogeneous domain.
Several studies have evaluated LLMs in medical contexts. Gao et al. (2023) assessed realistic medical tasks, though they did not include few-shot evaluations or modern LLM architectures. Kweon et al. (2024) and Adams et al. (2024) focused on realistic input documents like EHRs, without specifically evaluating biomedical LLMs. Chen et al. (2024) concentrated on biomedical models with MCQA tasks but did not explore practical tasks or compare performance before and after biomedical training. Liu et al. (2024) covered a broad range of tasks, including practical and biomedical models, though a systematic comparison of pre- and post-training performance was not conducted, and advanced models like Med42-v2 (Christophe et al., 2024a) and Meditron3 (OpenMeditron, 2024) were not considered.
Notably, none of these studies reported negative effects of biomedical training. In contrast, our study systematically compares $1212$ biomedical LLMs with general-domain counterparts on practical medical tasks and finds that biomedical training can lead to increased hallucinations and decreased accuracy, highlighting potential drawbacks of specialized training.
3 Evaluation Tasks
Our evaluation includes six clinical document tasks, encompassing information extraction, summarization, clinical reasoning, simplification, and coding. Two of these are introduced in this study to better assess the practical applications of LLMs in clinical settings. Table 1 summarizes the characteristics of these tasks. We provide prompt examples for each task in Figures 8, 9, 10, 11, 12, and 13 in Appendix B.3.
3.1 MeDiSumQA
MeDiSumQA is a novel medical open question-answering dataset we derived from MIMIC-IV (Johnson et al., 2021) discharge summaries.
3.1.1 Dataset Generation
Some of the discharge summaries contain a discharge letter to the patient. This letter summarizes the central information of the report in condensed and simple language and contains follow-up instructions. Based on these letters, we generate question-answer pairs using Meta’s Llama-3-70B-Instruct AI@Meta (2024). The questions are formulated from the patient’s view, and the answers are based on information that was given to the patient. For construction, we filter for documents containing a specific string111“You were admitted to the hospital” that marks the beginning of the discharge letter. The creation process is divided into three phases that are visualized in Figure 2.
In the first phase, we split the text into sentences (for details, see A.1). In the second phase, we prompted the LLM to generate question-answer pairs from the list of sentences for each document. We instruct the LLM to ground each answer in one or more sentences so we can check the quality manually. After completing the first two phases, we have a list of question-answer pairs. In the third phase, these candidates are manually reviewed by a medical expert, and high-quality examples are selected based on the following criteria:
  1. Detail: We observed that some information in discharge letters did not match the rest of the discharge summary due to medical details that were not communicated in the letter.
  2. Difficulty: We found that some generated answers were too obvious, or the questions were so specific that they gave away the answer.
  3. Ambiguity: Many questions lacked clear answers. For instance, a question about a procedure performed for diagnosis is answered by mentioning a CT scan. However, the full report also listed laboratory tests as diagnostically relevant.
Figure 7 in Appendix A.1 lists some examples of problematic pairs and selected pairs. As a last step, we remove the discharge letters from the original documents to pair the questions and answers with the resulting documents.
3.1.2 Task Qualities
This task requires models to perform multiple skills simultaneously. They must read and comprehend discharge summaries to fully understand the patient’s case, extract relevant details to answer specific questions about the patient’s condition and treatment and rephrase and simplify the information to ensure it is easily understandable by the patient. Additionally, the models need a robust understanding of medical knowledge and clinical guidelines to infer and provide appropriate follow-up advice accurately.
This multifaceted challenge assesses the LLM’s ability to integrate comprehension, extraction, communication, and domain-specific expertise in a clinical context.
3.2 MeDiSumCode
Figure 3: An overview of the MeDiSumCode dataset.
Coding of discharge summaries involves assigning International Classification of Diseases (ICD) (Organization, 2004) codes to diagnoses and procedures. This task is necessary for patient records, billing, and statistical analysis in healthcare.
3.2.1 Dataset Generation
Using MIMIC-IV, we generate an ICD-10 prediction dataset by leveraging the hosp module, which contains manually annotated ICD-10 codes for primary and secondary diagnoses and corresponding discharge summaries. Figure 3 shows that our process links discharge summaries to their respective ICD-10 codes using the unique hospital admission ID. This linkage allows us to create a cohesive dataset where each discharge summary is accurately paired with its diagnostic codes, providing both the inputs (discharge summaries) and the labels (ICD-10 codes) for model evaluation.
3.2.2 Task Qualities
This task is challenging for several reasons. First, it requires accurate identification of relevant diagnoses from complex and long clinical text. Second, the system must have detailed knowledge of the vast ICD-10 coding system, which contains over 70,000 codes, in order to accurately match these diagnoses. This requires a large database of codes and an understanding of the medical conditions they represent. Finally, matching diagnoses to the correct ICD-10 codes involves reasoning to combine both prior steps for a prediction.
MeDiSumCode assesses an LLM’s ability to extract relevant information from complex text accurately, possess detailed knowledge of a specialized coding system, and employ reasoning to integrate these elements for precise predictions.
3.3 MedNLI
The basis for MedNLI (Romanov and Shivade, 2018) are clinical notes from the MIMIC III (Johnson et al., 2016) dataset. Sentences are sampled from medical history sections and given to clinicians. Based on each sentence, generate three hypotheses—contradictory, neutral, and entailed—. Models are tested by predicting the logical relationship between the premise and these hypotheses.
The input length of this dataset is extremely small, making it well suited to measuring the ability for clinical reasoning regardless of the difficulty of LLMs on long inputs.
3.4 MeQSum
MeQSum (Ben Abacha and Demner-Fushman, 2019) consists of 1,00010001,0001 , 000 consumer health inquiries from the U.S. National Library of Medicine that medical experts manually summarized. This task provides a good complement to clinical documents written by medical staff, as it checks whether models are able to understand lay language.
MeQSum probes if LLMs are able to identify key information patient queries and reformulate them to brief, directed medically sound queries.
3.5 ProblemSummary
This task was first described by Gao et al. (2022) and utilizes clinical notes organized according to the widely recognized SOAP principle, which divides notes into Subjective, Objective, Assessment, and Plan sections (Weed, 1964). For this task, only the Subjective and Assessment sections were employed as inputs for the models, with the Objective portion being excluded. Based on this information, models are assigned to predict a patient’s current health problems.
Similar to MedNLI, the input length of this dataset is quite short, allowing us to test basic information extraction abilities without requiring a long context window.
3.6 LongHealth
The LongHealth (Adams et al., 2024) dataset contains 20 fictional patient records with various diseases. The reports are designed to be long and thus present an additional challenge to LLMs, as their performance has previously been shown to worsen with increasing input length (Levy et al., 2024). Evaluation is split into three sub-tasks: The first sub-task measures how well models can answer questions about long documents. Sub-task two increases the input length by including unrelated documents. The third sub-task removes the relevant document, and a new answer option is added to state that the information is not available.
This task measures a model’s comprehension of complex clinical documents, its ability to retain this performance on long inputs, and its tendency to hallucinate an answer when information is not available.
3.7 Metrics
For open-ended tasks, we report the F1-score between the model predictions and ground truth unigrams (ROUGE-1), bigram (ROUGE-2), and the longest common subsequence (ROUGE-L) [^2] https://huggingface.co/spaces/evaluate-metric/rouge (Lin, 2004). We compute the BERTScore (Zhang et al., 2019) on clinical documents to measure semantic similarity using an encoder trained on MIMIC III333emilyalsentzer/Bio_ClinicalBERT (Alsentzer et al., 2019). We first tuned the score rescaling baselines for MIMIC IV discharge summaries. For Problem Summaries and MeDiSumQA, we also extract the Unified Medical Language System (UMLS) (Bodenreider, 2004) entities with scispacy (Neumann et al., 2019) and compute their F1-score to consider medical abbreviations and synonyms. When evaluating MedDiSumCode, we calculate the ratio of valid ICD-10 codes. We use the python package icd10-cm [^4] https://pypi.org/project/icd10-cm/ to probe the validity of ICD-10 codes. We distinguish between exact match (EM) and the match of the first three characters of the codes, which is an approximate match (AP) based on the hierarchical structure of ICD-10 codes.
Following the SuperGLUE benchmark (Wang et al., 2019), we compute an aggregate score for both levels by averaging all over all datasets. For datasets with more than one metric, we first compute the average score of the dataset.
4 Experimental setup
Figure 4: The average level 1 and level 2 scores. Models that share the same base are displayed in the same color. Arrows indicate the performance difference to the base model after biomedical training.
We evaluated $2525$ language models, including biomedically trained models, their base models, and additional general-domain models as reference. Our evaluation aims to (1) measure the effects of continuous biomedical training, (2) assess whether biomedical models or general domain models are more suitable for specific medical scenarios, and (3) rank current openly available models.
4.1 Biomedical Models
Table 4 in B.2 shows the biomedical LLMs we evaluate in our experiments. Meditron-7B and 70B (Chen et al., 2023) are continuously pretrained from Llama-2-7B Touvron et al. (2023a). The pretraining dataset consists of 48.1B tokens (Chen et al., 2023). In contrast to Meditron, internistai/base-7b-v0.2 (Griot et al., 2024) was only trained on 2.3B tokens with a focus on data quality based on Mistral-7B-v0.1 (Jiang et al., 2023). BioMistral (Labrak et al., 2024) is another continuously pre-trained model that was based on Mistral-7B-Instruct-v0.1. They use weight-merging algorithms such as DARE (Yu et al., 2023) to combine the weights of their biomedical model and the original instruction-tuned model. A more recent model was published by Gururajan et al. (2024) and is based on Meta’s Llama3 (AI@Meta, 2024). They employed Direct Preference Optimization (DPO) (Rafailov et al., 2024) and supervised fine-tuning (SFT) to train instruction-tuned models. A similar fine-tuning procedure was used for Llama3-OpenBioLLM-8B and Llama3-OpenBioLLM-70B (Ankit Pal, 2024). However, no information on the training process has been provided yet. Med42-Llama3-8B and Med42-Llama3-70B (Christophe et al., 2024b) were trained on a mixture of various medical datasets, as well as general domain datasets using and iterative alignment approach (Christophe et al., 2024b). Finally, Meditron3-8B and Meditron3-70B (OpenMeditron, 2024) were trained on a similar dataset as the first Meditron. However, this generation was trained from instruction-tuned Llama3.1 models (Dubey et al., 2024) and the data also included instructions.
4.2 General-domain models
Apart from the previously mentioned models we also evaluated zephyr-7b-beta Tunstall et al. (2023), Mistral-7B-Instruct-v0.2 Jiang et al. (2023), Phi-3-mini-128k-instruct Abdin et al. (2024), Mixtral-8x7B and Mixtral-8x22B Jiang et al. (2024) to compare the performance of biomedical models with general-domain models.
4.3 Levels
Since the context length of Llama2-based models is insufficient to read all MIMIC discharge summaries and LongHealth documents, we split the evaluation into two levels: The first one with input sizes of less than 1.000 words on average and level two with more than 1.000 on average with long documents. Table 1 shows the average number of words per input for all tasks. We evaluate models with a context size of less than 8k tokens only on level 1 tasks.
4.4 Baselines
For each score, we provide a lower bound for meaningful predictions as a baseline to facilitate the interpretability of the scores. We use the expected value for random answer selection for the two data sets with finitely many answers, MedNLI and LongHealth. For MeQSum and MeDiSumQA, we use the unchanged input as a prediction. In the case of Problem Summary, we extract all UMLS entities from the input and use them as predictions. Finally, for MeDiSumCode, we collect all ICD-10 codes occurring in the dataset and draw random codes from this set.
5 Results
We present the benchmark results in Figure 4, with detailed scores for all models available in Table 5 in Appendix C.
Figure 4 illustrates that models often perform lower after biomedical training, especially when the respective base model performs well. Notably, this applies to all models based on Llama3 and Llama3.1 except Llama3-Med42-8B which performs better on simple level 1 tasks, but worse on level 2 tasks.
Models based on Mistral-7B, which is the lowest-scoring general-domain model, show a performance improvement. However, the improvement of BioMistral-7B is only marginal, and internistai/base-7b-v0.2 only improves on level 2 tasks. Additionally, they do not demonstrate a clear advantage over zephyr-7b-beta, another general-domain model based on Mistral-7B.
Generally, Llama3 and Llama3.1-based biomedical models show a clearer performance drop on level 2 tasks than level 1 tasks, indicating the expressiveness of LongHealth and the newly introduced tasks, MeDiSumQA and MeDiSumCode.
LH Task3 Valid Codes
BioMistral-7B +4.15 +17.26
BioMistral-7B-DARE +0.95 +18.79
internistai/base-7b-v0.2 +45.55 +16.32
Llama3-OpenBioLLM-8B -40.05 -10.77
Llama3-Med42-8B -12.7 -6.8
Llama3-Aloe-8B-Alpha -22.55 -17.09
Llama3-OpenBioLLM-70B -28.80 -20.29
Llama3-Med42-70B -79.15 -15.39
Meditron3-8B -52.15 -49.19
Meditron3-70B -54.6 -4.76
Table 2: Mistral-7B-v0.1 , Meta-Llama-3-(8B/70B) and Meta-Llama-3.1-(8B/70B) based models on LongHealth task 3 and percentage of valid ICD-10 codes in MeDiSumCode
5.1 Error Analysis
Major contributions to the performance drop of biomedical models are task 3 of LongHealth and the MeDiSumCode valid code scores. Table 2 shows the performance difference between biomedical and their base models on these metrics. All models based on Mistral-7B show a clear performance improvement, while all Llama3-based models perform worse.
Since LongHealth task 3 measures how often a model returns no answer if the requested information is not in the input document, it indicates the frequency of hallucinations. Similarly, the valid code scores in MeDiSumCode show models often invent ICD-10 codes. Models that have a low valid code score often start to count up the numbers in ICD codes instead of predicting valid codes. Examples of this behavior are listed in Appendix C.1. LongHealth task 3 reveals a drastic change from Meta-Llama-3-8B-Instruct, which scored $56.2556.25$, to Llama3-OpenBioLLM-8B, which only scored $1.551.55$. Similarly, Llama3-OpenBioLLM-70B also performs poorly compared to Meta-Llama-3-70B-Instruct on this task.
In addition to increased hallucinations, we observed that biomedically fine-tuned models often exhibit generation loops, wherein they repetitively produce the same sequences without reaching a conclusion. This looping behavior significantly hampers the models’ ability to generate coherent and relevant responses. Examples of this phenomenon have been observed in various tasks, highlighting the consistent nature of these generation loops.
Another common issue identified in poorly performing models is their inability to adhere to the prescribed instruction format, particularly in tasks involving long inputs at level 2. This problem is notably pronounced in tasks such as LongHealth, where the expectation is for models to produce answers in a strict, parseable format.
6 Discussion
This section explores the implications of our findings in comparison to existing evaluations, and the broader effects of biomedical training on model performance.
Model MCQA Level 1 Level 2
MEDITRON-7B +6.07 -7.08 -
MEDITRON-70B +3.63 -4.59 -
BioMistral-7B +4.13 +0.26 +0.71
BioMistral-7B-DARE +4.57 +2.93 +2.7
internistai/base-7b-v0.2 - -2.07 +5.52
Llama3-OpenBioLLM-8B -0.63 -15.17 -13.54
Llama3-OpenBioLLM-70B +1.46 -4.78 -10.45
Llama3-Med42-8B +0.47 +2.51 -1.4
Llama3-Med42-70B +2.8 -7.57 -15.14
Llama3-Aloe-8B-Alpha +2.21 -5.87 -8.67
Meditron3-8B - -2.76 –15.04
Meditron3-70B - -2.18 -8.51
Table 3: Reported performance difference after biomedical pretraining on MCQA tasks and our evaluation for models based on Llama-2-(7B/70B), Mistral-7B-v0.1, Meta-Llama-3-(8B/70B)-Instruct and Meta-Llama-3.1-(8B/70B)-Instruct.
6.1 Discrepancy to previous benchmarks
[…]
However, it is important to note that MCQA tasks primarily assess the medical knowledge of models without accounting for the challenges encountered in clinical practice, such as medical jargon, abbreviations, typos, irregular formatting, incomplete information, strict task formats, patient-friendly language, and long inputs. In contrast, our analysis, reveals a notable decline in performance after biomedical fine-tuning for many models. This decline suggests that our evaluation tasks, being more general and reflective of real-world scenarios, may expose deficiencies in the models’ retained general-domain capabilities following their specialization in biomedical tasks.
Notably, models based on the Mistral-7B architecture present mixed results, with some showing improvements while others experience performance deterioration. Despite these exceptions, the overall trend indicates a more pronounced decline in task performance than the improvements seen on MCQA benchmarks. This pattern suggests that biomedical fine-tuning may have an overall negative impact on the models’ ability to perform general clinical tasks, revealing a trade-off between domain-specific optimization and broader general-domain task performance.
6.2 Effect of biomedical training
Our analysis reveals a notable decline in performance for several models following biomedical training, observable after both continuous pre-training and methods like SFT and DPO. An outlier in this trend is BioMistral-DARE, which employed weight merging with the original instruction-tuned checkpoint. This suggests that such a training approach may mitigate the adverse effects of fine-tuning.
However, the superior performance of Mistral-7B-Instruct-v0.2, which shares the same architecture as BioMistral-DARE, implies that an enhanced general domain pre-training dataset might have a more profound impact than fine-tuning alone.
Additionally, many models trained using SFT relied on generated data, hinting that performance issues may stem from data quality. internistai/base-7b-v0.2, which was trained on high-quality data, showed the highest improvement in performance on level 2 tasks, supporting the hypothesis that data quality is crucial.
All biomedical models that showed improvement are based on the relatively low-scoring Mistral-7B-Instruct-v0.1. This leads us to our final hypothesis: these improvements might address gaps that more recent general-domain models, such as Llama-3 and Mistral-7B-Instruct-v0.2, have already overcome. This is supported by Tables 2 and 3 showing improvements of biomedical Mistral-7B models and performance deterioration of biomedical Llama3 models.
Therefore, we conclude that biomedical training with currently existing methods and datasets is not beneficial for models in clinical settings. Future research could be focused on novel training methods that counteract model performance deterioration or focus on datasets that address this issue.
7 Conclusion
Our study reveals that biomedical LLMs are not competing effectively with general-domain models in practical medical settings. While some biomedical models have shown marginal improvements, many recent models are underperforming. Fine-tuning these models with domain-specific data often leads to reduced performance, introducing hallucinations and decreased model stability. This stands in contrast to traditional MCQA evaluations, where biomedical models have previously demonstrated superior performance. Our evaluation provides a more practical and accurate assessment of LLM capabilities in real-world healthcare settings. To support further progress in this field, we are open-sourcing our evaluation scripts, allowing for broader validation and replication of these results. We believe future research should focus on addressing the highlighted issues to enhance the effective deployment of LLMs in medical practice.
Limitations
Our study has several limitations that should be considered. Due to the significant computational resources required to run LLMs with up to 141 billion parameters, we did not explore the impact of various model configurations, such as temperature settings, or advanced techniques like chain-of-thought prompting on model performance. Future research should investigate these aspects to gain a more comprehensive understanding of their effects. Additionally, our new datasets are based on publicly available data from MIMIC-IV. As such, we cannot completely prevent data contamination. This limitation underscores the need for future research into robust methods for mitigating data contamination, which is crucial for ensuring the validity of any public LLM benchmark. While we presented novel insights in this paper, their application to clinical data requires further investigation. Future work should refine these methods to enhance their applicability and reliability in clinical settings. Furthermore, our evaluation primarily focused on tasks involving clinical documents and their relevance, but it was not conducted in a realistic clinical setting. Therefore, extensive evaluation through prospective clinical trials is necessary to meet the required safety levels before applying these models to clinical environments.
Appendix A Task Details
A.1 MeDiSumQA Generation
Figure 5: The prompt we use to split the discharge instructions into sentences.
Figure 6: The prompts we are using to generate MeDiSumQA.
We do not use a classic sentence splitter (e.g., NLTK), as these do not work reliably due to the irregular formatting and placeholders introduced by anonymization. Instead, we prompt an LLM to split the text into sentences without changing anything else. We make sure that the LLM has not changed the sentences by checking whether the sentences can be found in the original document via exact string matching. Figure 5 shows our prompt for this.
Figure 7: This figure shows representative examples of issues we detected with some of the generated questions-answer pairs.
We manually select the final examples for the dataset based on correctness, difficulty, and ambiguity. Figure 7 shows representative examples of pairs we filtered out.
Model Name Base Model Type of Training
Meditron-7B llama2-7b Continued pretraining
Internist.ai 7B Mistral-7B-v0.1 Continued pretraining + SFT
BioMistral-7B Mistral-7B-Instruct-v0.1 Continued pretraining
BioMistral-7B-DARE Mistral-7B-Instruct-v0.1 Continued pretraining +DARE
Llama3-OpenBioLLM-8B Meta-Llama-3-8B-Instruct SFT + DPO
Llama3-Med42-8B Meta-Llama-3-8B-Instruct SFT + DPO
Llama3-Aloe-8B-Alpha Meta-Llama-3-8B-Instruct SFT + DPO
Meditron3-8B Meta-Llama-3.1-8B-Instruct -
Meditron-70B Llama-2-70b Continued pretraining
Llama3-OpenBioLLM-70B Meta-Llama-3-70B-Instruct SFT + DPO
Llama3-Med42-70B Meta-Llama-3-8B-Instruct SFT + DPO
Meditron3-70B Meta-Llama-3.1-70B-Instruct -
Table 4: Evaluated Biomedical Models
Appendix B Experimental setup
B.1 Computational Resources
All experiments were conducted on an NVIDIA DGX A100 640GB node with 8x NVIDIA A100 80GB Tensor Core GPUs within three days, resulting in approximately 1536 GPU hours.
B.2 Models
Table 4 lists all biomedical models we evaluated.
B.3 Prompting
We apply few-shot prompting and use the instruction template on Hugging Face for the instruction-tuned models. For the other models, we concatenate the system prompt, few-shot examples, and user prompt into one string separated by double newlines. For the level one evaluation, we performed 3-shot prompting. For level two, we provide one shot with the exception of LongHealth, where we provide no examples due to the content length.
Figures 8, 9, 10, 11, 12, and 13 are showing the prompt formats we are using for the different benchmark tasks. If the input length allowed this, we also included sample texts from the datasets.
Figure 8: MeQSum prompt format with example.
Figure 9: Problem Summary prompt format with example.
Figure 10: MedNLI prompt format with example.
Figure 11: LongHealth prompt format.
Figure 12: MeDiSumQA prompt format.
Figure 13: MeDiSumCode prompt format.
Appendix C Results
Table 5 shows the detailed benchmark results for all models.
Figure 14: Biomedical models that show the described counting behavior compared to their base model.
C.1 Error Analysis
Figure 14 shows some examples of the described type of error with regard to counting.
MedNLI Problem Summary MeQSum
Level 1 Score Acc R-L R-1 R-2 BERT F1 UMLS F1 R-L R-1 R-2 BERT F1
baseline 25.13 33.33 6.22 8.53 1.58 61.07 8.24 14.96 18.99 7.21 58.62
Llama-2-7b 20.11 29.51 5.97 7.35 2.11 59.45 9.06 7.16 8.58 2.89 37.5
Meditron-7B 13.03 2.39 11.29 13.43 4.81 63.35 15.21 6.83 7.90 2.26 43.39
Llama-2-70b 35.17 76.27 7.13 8.77 3.15 59.61 14.34 3.75 4.15 1.05 33.59
Meditron-70B 30.58 63.52 7.22 8.91 3.21 60.08 13.91 2.95 3.24 0.76 31.27
Mistral-7B-v0.1 33.99 67.54 7.07 8.99 3.17 60.74 15.67 6.28 7.86 2.49 44.6
Mistral-7B-Instruct-v0.1 40.31 64.79 14.46 18.6 6.26 65.79 20.08 21.85 25.26 11.17 66.15
zephyr-7b-beta 42.69 68.45 14.81 19.91 5.98 67.39 19.2 25.66 29.81 12.33 68.85
BioMistral-7B 40.57 62.75 16.9 20.89 8.4 59.03 20.12 25.89 28.46 13.31 67.93
BioMistral-7B-DARE 43.24 66.76 18.81 22.92 8.93 68.93 22.65 26.16 29.69 13.49 68.68
internistai/base-7b-v0.2 38.24 76.34 13.11 16.79 5.63 62.94 17.22 7.65 9.54 3.44 40.34
Phi-3-mini-128k-instruct 41.70 57.25 19.8 23.72 8.47 70.27 21.06 32.05 35.85 15.81 73.01
Mistral-7B-Instruct-v0.2 46.45 69.93 19.57 25.59 8.92 69.64 22.07 33.54 37.47 16.61 73.47
Mixtral-8x7B-Instruct-v0.1 47.81 76.48 17.44 23.39 7.7 68.56 19.51 32.47 36.38 16.86 72.8
Mixtral-8x22B-Instruct-v0.1 51.88 84.93 17.97 22.6 8.17 69.29 22.31 36.15 39.94 19.45 75.02
Meta-Llama-3-8B-Instruct 48.37 74.08 22.7 28.52 9.87 71.45 25.32 32.2 36.49 16.37 72.74
Llama3-OpenBioLLM-8B 33.20 44.93 10.82 13.62 4.03 64.14 15.67 26.21 29.41 14.03 62.39
Llama3-Med42-8B 50.88 77.46 24.43 29.18 10.76 72.51 25.12 36.03 40.32 19.83 74.93
Llama3-Aloe-8B-Alpha 42.50 73.94 9.47 12.3 4.06 65.56 15 23.78 27.02 12.31 66.01
Meta-Llama-3.1-8B-Instruct 50.32 79.08 20 26.14 8.53 70.79 23.5 35.69 39.55 18.11 75.03
Meditron3-8B 47.56 74.01 18 22.39 8.1 69.26 21.8 33.61 37.6 17.84 74.01
Meta-Llama-3-70B-Instruct 52.36 79.37 25.43 33.16 13.01 73 29.12 36.57 40.2 19.3 75.74
Llama3-OpenBioLLM-70B 47.57 80.85 12.1 16.67 5.58 66.51 17.74 30.72 34.31 15.55 71.99
Llama3-Med42-70B 44.79 76.13 14.12 18.68 5.52 65.76 17.59 25.69 29.41 12.56 67.98
Meta-Llama-3.1-70B-Instruct 54.37 84.86 25.26 32.62 12.63 72.76 29.43 37.62 41.23 19.68 76.36
Meditron3-70B 52.19 82.61 22.72 27.69 11.11 71.43 26.22 35.42 39.09 19.08 74.91
LongHealth MeDiSumQA MeDiSumCode
Level 2 Score Task 1 Task 2 Task 3 R-L R-1 R-2 BERT F1 UMLS F1 EM F1 AP F1 Valid Code
baseline 24.86 20.00 20.00 16.66 13.11 15.76 2.82 60.22 12.74 0.88 3.44 100.00
Mistral-7B-v0.1 9.55 0.60 0.55 0.15 5.93 7.16 1.59 53.47 7.47 0.77 5.32 33.21
Mistral-7B-Instruct-v0.1 23.12 45.75 40.65 3.65 16.62 21.34 6.95 65.68 16.77 0.57 3.78 37.25
zephyr-7b-beta 28.17 42.9 30.5 26.35 13.02 17.68 4.98 64.08 13.92 2.31 12.00 71.27
BioMistral-7B 23.83 38.05 34.25 7.8 14.65 17.81 5.46 59.01 16.88 1.67 9.92 54.51
BioMistral-7B-DARE 25.82 46.00 40.85 4.6 17.01 20.87 6.92 65.17 18.45 1.20 6.66 56.04
internistai/base-7b-v0.2 28.64 52.75 30.6 49.2 9.24 11.94 3.23 61.82 13.00 1.87 10.25 53.57
Phi-3-mini-128k-instruct 24.39 27.25 23.9 0.00 18.47 22.45 6.94 65.72 16.52 0.43 3.63 86.24
Mistral-7B-Instruct-v0.2 38.94 67.2 62.4 42.45 21.8 27.47 9.19 68.43 20.3 3.08 18.23 68.76
Mixtral-8x7B-Instruct-v0.1 42.57 76.5 73.65 24.2 20.72 26.4 8.96 67.74 20.25 10.49 28.99 82.87
Mixtral-8x22B-Instruct-v0.1 51.23 79.00 73.90 86.30 21.81 27.75 9.42 68.51 22.70 15.95 35.96 79.84
Meta-Llama-3-8B-Instruct 38.98 68.3 66.55 41.6 22.44 28.15 9.63 68.62 22.74 3.95 17.55 61.93
Llama3-OpenBioLLM-8B 25.44 37.55 41.75 1.55 22.89 27.95 10.45 68.7 22.15 0.84 4.84 51.16
Llama3-Med42-8B 37.58 74.35 70.3 28.9 21.86 26.81 9.29 68.15 22.23 4.32 16.22 55.13
Llama3-Aloe-8B-Alpha 30.31 66.75 63.3 19.05 11.03 14.97 4.57 63.87 12.68 1.77 12.78 44.84
Meta-Llama-3.1-8B-Instruct 45.61 76.65 75.4 59.45 26.23 31.52 11.45 70.38 24.82 6.72 22.54 71.08
Meditron3-8B 30.57 73.65 70.5 7.3 23.52 28.71 10.61 69.1 23.63 2.23 6.25 21.89
Meta-Llama-3-70B-Instruct 56.00 81.65 77.90 91.70 26.2 32.5 11.93 70.24 25.78 19.65 39.2 93.94
Llama3-OpenBioLLM-70B 45.56 80.2 75.60 62.90 21.85 27.8 9.54 68.43 22.47 7.37 20.24 73.65
Llama3-Med42-70B 40.86 79.05 77.55 12.55 14.55 19.72 5.53 65.53 15.82 14.2 33.18 78.55
Meta-Llama-3.1-70B-Instruct 57.68 84.95 79.65 98.5 25.18 30.79 11.2 70.09 25.6 21.72 40.62 95.95
Meditron3-70B 49.17 81.55 77.55 43.9 24.64 29.93 11.1 69.66 25.21 16.68 35.35 91.19
Table 5: The benchmark results for level 1 and level 2. Models with the same base model are grouped. †Indicates biomedical training. The relative level of the scores for each column is colored from red for low scores to green for high scores.